# External Attention Transformer: A Robust AI Model for Identifying Initial Eccentricity Signatures in Binary Black Hole Events in Simulated Advanced LIGO Data

### v1

@@ 1 Introduction
It turns out that initial orbital eccentricities, component spins, masses, and eccentricities of compact binaries provide crucial clues for distinguishing between different formation channels, as they encapsulate information about the formation history of the system [6, 7]. As the eccentricity of a binary system rapidly decreases due to the emission of gravitational wave radiation [8], isolated binary black holes that inspiral in eccentric orbits are expected to have negligible eccentricities by the time they reach the sensitivity range of LVK observations. GW emissions from binaries formed through the isolated evolution process are expected to be detected as they inspiral on quasi-circular orbits. In this scenario, two stars evolve around each other with minimal interaction with their environment, eventually forming compact objects and merging within the age of the Universe [9, 6, 7]. However, in the case of dynamical formation, certain processes can cause the detected eccentricity to be non-negligible. Dynamical formation occurs in populated environments, such as globular clusters, young star clusters, and galactic nuclei [10, 11, 12]. Merger progenitors formed through the dynamical formation channel may be observed with eccentricities $e_{10}$ $>$ 0.1 in the LVK sensitive frequency range (above 10 Hz) [13, 14, 15, 16, 6, 17].
@@ 2.2 Preprocessing: creating the Synthetic Data
We employ the SEOBNRv5EHM waveform model [53, 54] from the effective-one-body numerical-relativity (EOBNR) waveform model family [55] using the pyseobnr package [56] and LALSuite [57, 58] to generate simulated eccentric gravitational wave (GW) waveforms. This model represents an order of improvement over the SEOBNRv4EHM model described in Ref. ([59], [60]) and includes higher-order multipoles: $(\ell,|m|)=(2,2),(2,1),(3,2),(3,3),(4,3),(4,4)$. For this work, we focus on the dominant $(\ell,|m|)=(2,2)$ mode in our simulations. The chirp mass $M_{c}$ of each binary is drawn uniformly from the range (15, 60) $M_{\odot}$. The mass ratio $q=m_{2}/m_{1}$ spans the range 0.25 to 1. The initial orbital phase $\zeta$ is sampled uniformly over $(0,2\pi)$, and the eccentricity at 15 Hz, $e_{15,\mathrm{Hz}}$, is drawn uniformly from $(0.001,0.35)$. Distances are randomly sampled within the range 100 to 1500 Mpc. The sky position (right ascension and declination) and orientation (inclination and polarization angles) are each uniformly sampled over the sky and the polarization sphere, respectively. For simplicity, we neglect the spins of the binary black hole components. Waveforms are generated at a sample rate of 4096 Hz, with a frequency resolution of 1/16 Hz. Following the methodology outlined in Ref. [61], we generate spectrograms of the simulated signals using PyCBC [62]. Each waveform is 30 seconds long, spanning the time window $[-15,15]$ seconds, and is injected into 40 seconds of Gaussian noise ($[-20,20]$ seconds) using the O4 sensitivity curve for the LIGO detectors [^3] https://dcc.ligo.org/ligo-t2000012/public. The merger time is set to zero for all synthetic signals. To reduce computational costs, we consider only the data in the $[-10,2]$ seconds window for generating spectrograms. We have checked that this time range is sufficient to capture the full signal duration, even for the longest signals in our simulations, which correspond to $M_{c}$=15 $M_{\odot}$ and q=1. For each simulated dataset, we calculate the matched-filter signal-to-noise ratio (SNR) and retain only those samples with LIGO Hanford–LIGO Livingston network SNR greater than 10. The resulting signals are band-pass filtered in the $[20,512]$ Hz frequency range and whitened, after which the Q-transform is applied to the whitened signal. The resulting Q-transform spectrograms are resized to $256\times 256$ pixels and stacked to construct the dataset for subsequent visual processing tasks.
@@ 3 Implications of Our Approach
To evaluate the performance of our model, we generate several independent test datasets, each with fixed values of chirp mass and eccentricity, across different network SNR ranges: (10–15), (15–20), (20–30), (30–40), (40–50), and (50–60). For each network SNR range, we generate 500 events. On average, the model required on the order of 100 ms to process a single event, as determined from 100 repeated inferences. Figure 3 shows the recovered distributions of $e_{15,\mathrm{Hz}}$ = 0.07 and $M_{c}$ = 40 $M_{\odot}$ for 500 events in each network SNR bin. In each subpanel, the parameters are recovered within 1 $\sigma$, and the size of the contours decreases as the network SNR increases, indicating improved precision. Figure 4 presents a similar analysis for a higher eccentricity value of 0.2.
@@ 4 Conclusions
A significant advantage of our method is its rapid and computationally efficient training process, enabled by the external attention mechanism, which employs external memory to capture general patterns effectively across subsamples. Specifically, we trained the model on 24,000 gravitational wave spectrograms, corresponding to the training dataset, using a single GPU (NVIDIA GeForce RTX 2070 with 8 GB VRAM). During training, 20% of the training data was used for validation, and the entire training process was completed in less than one hour. The average inference time per event, measured over 100 runs, was on the order of 100 ms. Testing across diverse datasets, as discussed in Section 3, demonstrates that our model achieves sufficient precision in determining eccentricity. Recognizing the intrinsic degeneracy between eccentricity and chirp mass, our approach simultaneously estimates both parameters to improve robustness and reliability. Furthermore, the model is trained on data simulating the current LVK O4 observational run, making it immediately applicable to ongoing analyses. Due to its efficient training capability, this model can easily be retrained or updated for future LVK observational runs, highlighting its practical utility for rapid and accurate GW parameter estimation. These considerations imply that our approach should allow us to provide independent constraints on the reported initial eccentricities of a few O3 GW events [18, 25, 26]. This is currently under investigation.

### v2

@@ 1 Introduction
It turns out that initial orbital eccentricities, component spins, masses, and eccentricities of compact binaries provide crucial clues for distinguishing between different formation channels, as they encapsulate information about the formation history of the system [6, 7]. As the eccentricity of a binary system rapidly decreases due to the emission of gravitational wave radiation [8], isolated binary black holes that inspiral in eccentric orbits are expected to have negligible eccentricities by the time they reach the sensitivity range of LVK observations. GW emissions from binaries formed through the isolated evolution process are expected to be detected as they inspiral on quasi-circular orbits. In this scenario, two stars evolve around each other with minimal interaction with their environment, eventually forming compact objects and merging within the age of the Universe [9, 6, 7]. However, in the case of dynamical formation, certain processes can cause the detected eccentricity to be non-negligible. Dynamical formation occurs in populated environments, such as globular clusters, young star clusters, and galactic nuclei [10, 11, 12]. Merger progenitors formed through the dynamical formation channel may be observed with eccentricities $e_{10}$ >> 0.1 in the LVK sensitive frequency range (above 10 Hz) [13, 14, 15, 16, 6, 17].
@@ 2.2 Preprocessing: creating the Synthetic Data
We employ the SEOBNRv5EHM waveform model [53, 54] from the effective-one-body numerical-relativity (EOBNR) waveform model family [55] using the pyseobnr package [56] and LALSuite [57, 58] to generate simulated eccentric gravitational wave (GW) waveforms. This model represents an order of improvement over the SEOBNRv4EHM model described in Ref. ([59], [60]) and includes higher-order multipoles: $(\ell,|m|)=(2,2),(2,1),(3,2),(3,3),(4,3),(4,4)$. For this work, we focus on the dominant $(\ell,|m|)=(2,2)$ mode in our simulations. The chirp mass $M_{c}$ of each binary is drawn uniformly from the range (15, 60) $M_{\odot}$. The mass ratio $q=m_{2}/m_{1}$ spans the range 0.25 to 1. The initial orbital phase $\zeta$ is sampled uniformly over $(0,2\pi)$, and the eccentricity at 15 Hz, $e_{15,\mathrm{Hz}}$, is drawn uniformly from $(0.001,0.35)$. Distances are randomly sampled within the range 100 to 1500 Mpc. The sky position (right ascension and declination) and orientation (inclination and polarization angles) are each uniformly sampled over the sky and the polarization sphere, respectively. For simplicity, we neglect the spins of the binary black hole components. Waveforms are generated at a sample rate of 4096 Hz, with a frequency resolution of 1/16 Hz. Following the methodology outlined in Ref. [61], we generate spectrograms of the simulated signals using PyCBC [62]. Each waveform is 30 seconds long, spanning the time window [15,15][-15,15] seconds, and is injected into 40 seconds of Gaussian noise ([20,20][-20,20] seconds) using the O4 sensitivity curve for the LIGO detectors [^3] https://dcc.ligo.org/ligo-t2000012/public. The merger time is set to zero for all synthetic signals. To reduce computational costs, we consider only the data in the [10,2][-10,2] seconds window for generating spectrograms. We have checked that this time range is sufficient to capture the full signal duration, even for the longest signals in our simulations, which correspond to $M_{c}$=15 $M_{\odot}$ and q=1. For each simulated dataset, we calculate the matched-filter signal-to-noise ratio (SNR) and retain only those samples with LIGO Hanford–LIGO Livingston network SNR greater than 10. The resulting signals are band-pass filtered in the $[20,512]$ Hz frequency range and whitened, after which the Q-transform is applied to the whitened signal. The resulting Q-transform spectrograms are resized to $256\times 256$ pixels and stacked to construct the dataset for subsequent visual processing tasks.
@@ 3 Implications of Our Approach
To evaluate the performance of our model, we generate several independent test datasets, each with fixed values of chirp mass and eccentricity, across different network SNR ranges: (10–15), (15–20), (20–30), (30–40), (40–50), and (50–60). For each network SNR range, we generate 300 events. On average, the model required on the order of 100 ms to process a single event, as determined from 100 repeated inferences. We trained our model independently 10 times, and for each run, we saved the best-performing model based on the minimum validation loss, following the procedure described in subsection 2.3. Figure 3 shows the recovered distributions of $e_{15,\mathrm{Hz}}$ = 0.07 and $M_{c}$ = 40 $M_{\odot}$ for 300 events in each network SNR bin. In each subpanel, the parameters are recovered within 1 $\sigma$, and the size of the contours decreases as the network SNR increases, indicating improved precision. Figure 4 presents a similar analysis for a higher eccentricity value of 0.2.
@@
[…]
To better characterize prediction confidence in Figures 3-6, we evaluate model uncertainty using a deep-ensemble approach. We performed inference using all 10 trained models on several datasets and computed the standard deviation of the predictions for each event. We find that when the SNR is low, the main issue is the strong noise in the data. This noise makes it difficult for the model to identify true patterns, leading to greater variation in its predictions. In other words, most of the uncertainty at low SNR is due to the data, not limitations of the model. As the SNR increases and the signal becomes more prominent relative to the noise, the model’s predictions become much more precise and confident. In this high SNR regime, model uncertainty is comparable to the total uncertainty. The ensemble shows no significant improvement over a single model.
3.1 Quantitative Performance Metrics
To further quantify model performance, we report the mean absolute error (MAE), root mean squared error (RMSE), and their associated 95% confidence intervals across bins of $e_{15,\mathrm{Hz}}$ and $M_{c}$, separately for low and high SNR categories. For this evaluation, we generated a new test dataset, with the data binned by $e_{15,\mathrm{Hz}}$, $M_{c}$ , and SNR using the following bin edges: $e_{15,\mathrm{Hz}}$ bins of (0.05–0.10), (0.10–0.15), (0.15–0.20), (0.20–0.25), and (0.25–0.30); $M_{c}$ bins of (15–25), (25–35), (35–45), and (45–55) $M_{\odot}$; and SNR bins of (10–15), (15–20), (20–30), (30–40), (40–50), and (50–60). We used $N=100$ test samples in each $(e_{15\,\mathrm{Hz}},\,M_{c},\,\mathrm{SNR})$ bin combination. Throughout this analysis, we define low SNR as $10<SNR<20$ and high SNR as $20<SNR<60$. The detailed results are summarized in Table 1 for $e_{15,\mathrm{Hz}}$ bins and Table 2 for $M_{c}$ bins. These quantitative metrics provide a rigorous assessment of the model’s predictive accuracy across the relevant parameter space. As expected, the model achieves better performance for lower $M_{c}$ systems, whereas the errors increase for higher $M_{c}$. This trend, discussed in Section 2.3, arises because the time-frequency maps become shorter for higher $M_{c}$, leading to greater uncertainties in the recovered $M_{c}$. For $e_{15,\mathrm{Hz}}$ the model shows consistently reliable performance across the explored range.
Table 1: MAE and RMSE for $e_{15,\mathrm{Hz}}$ bins with 95% confidence intervals over ten independent runs.
[…]
0.05–0.10 $0490^{+0.0027}_{-0.0032} 0.0$$637^{+0.0033}_{-0.0029} 0.0$$216^{+0.0009}_{-0.0010} 0.0$$0281^{+0.0015}_{-0.0012}$
0.10–0.15 $0424^{+0.0022}_{-0.0021} 0.0$$529^{+0.0025}_{-0.0026} 0.0$$244^{+0.0011}_{-0.0011} 0.0$$0316^{+0.0015}_{-0.0013}$
0.15–0.20 $0422^{+0.0020}_{-0.0020} 0.0$$519^{+0.0022}_{-0.0024} 0.0$$247^{+0.0012}_{-0.0011} 0.0$$0323^{+0.0017}_{-0.0016}$
0.20–0.25 $0421^{+0.0023}_{-0.0023} 0.0$$531^{+0.0026}_{-0.0025} 0.0$$206^{+0.0009}_{-0.0010} 0.0$$0265^{+0.0013}_{-0.0013}$
0.25–0.30 $0415^{+0.0023}_{-0.0025} 0.0$$535^{+0.0027}_{-0.0032} 0.0$$181^{+0.0011}_{-0.0008} 0.0$$0251^{+0.0018}_{-0.0019}$
Table 2: MAE and RMSE for $\mathcal{M}_{c}$ bins with 95% confidence intervals over ten independent runs.
[…]
15–25 $0.8002^{+0.0500}_{-0.0460}$ $1.1008^{+0.1333}_{-0.0995}$ $0.4813^{+0.0199}_{-0.0184}$ $0.6084^{+0.0208}_{-0.0210}$
25–35 $1.3042^{+0.0871}_{-0.0776}$ $1.8262^{+0.1346}_{-0.1316}$ $0.6030^{+0.0257}_{-0.0268}$ $0.8116^{+0.0636}_{-0.0556}$
35–45 $2.4309^{+0.1292}_{-0.1289}$ $3.2095^{+0.1688}_{-0.1703}$ $1.1561^{+0.0607}_{-0.0655}$ $1.6413^{+0.0921}_{-0.0912}$
45–55 $2.8756^{+0.1432}_{-0.1326}$ $3.6514^{+0.1905}_{-0.1673}$ $1.6351^{+0.0773}_{-0.0662}$ $2.1609^{+0.1087}_{-0.0938}$
@@ 4 Conclusions
A significant advantage of our method is its rapid and computationally efficient training process, enabled by the external attention mechanism, which employs external memory to capture general patterns effectively across subsamples. Specifically, we trained the model on 24,000 gravitational wave spectrograms, corresponding to the training dataset, using a single GPU (NVIDIA GeForce RTX 2070 with 8 GB VRAM). During training, 20% of the training data was used for validation, and the entire training process was completed in less than one hour. The average inference time per event, measured over 100 runs, was on the order of 100 ms. We implemented ensemble learning to quantify model uncertainty. Testing across diverse datasets, as discussed in Section 3, demonstrates that our model achieves sufficient precision in determining eccentricity. Our findings show that the ensemble method does not improve performance compared to the single model. Recognizing the intrinsic degeneracy between eccentricity and chirp mass, our approach simultaneously estimates both parameters to improve robustness and reliability. Furthermore, the model is trained on data simulating the current LVK O4 observational run, making it immediately applicable to ongoing analyses. Due to its efficient training capability, this model can easily be retrained or updated for future LVK observational runs, highlighting its practical utility for rapid and accurate GW parameter estimation. These considerations imply that our approach should allow us to provide independent constraints on the reported initial eccentricities of a few O3 GW events [18, 25, 26]. This is currently under investigation.